BSM Berry is now running fully automatic. It is handling the northern survey targets as well as V339 Del, X Cyg and XZ Cet.
BSM_Hamren is essentially operational, when Bob Stine is in residence. It needs a slight modification on the image transfer operation, and a replaced focuser controller (which is a 1-minute swap).
BSM_NM burned out its power supply; Bill Stein has a replacement and will install it next week. We should be back in operation at that time. We are also installing the web power strip so that hardware can be turned on/off remotely.
BSM_South is fully operational with no problems. Way to go, Peter!
W30's mount controller was getting intermittent. Phil Coker has donated his 30cm Meade to replace W30. This will be called COKER30 in his honor. It is the ACF model, so the image quality may improve. Bill Goff is currently finalizing installation and checkout; data should start flowing soon.
SRO is fully operational with no problems. John knows his stuff!
OC61 is fully operational with no problems.
TMO61 is fully operational with no problems. Jon Holtzman just installed a Boltwood cloud sensor so that we have more remote information for open/close.